OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. (2 points) Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. (2 points) f(x) = 4/x and g(x) = 4/x

OpenStudy (anonymous):

I think it might be f(g(x)) = f(4/x) = 8/x^2 = x f(f(x)) = g(4/x) = 8/x^2 = x but I am not sure.

geerky42 (geerky42):

So plug in f(x) into x from g(x): \(g( f(x) )\), which will get you from \(g(x) = \dfrac{4}{x}\) to \(g(f(x)) = \dfrac{4}{~\dfrac{4}{x}~}\)

OpenStudy (anonymous):

So the answer is 4/4/x = x

OpenStudy (anonymous):

Or no?

geerky42 (geerky42):

You could show that \(\dfrac{4}{~\dfrac{4}{x}~} = \dfrac{4x}{4}=x\). Do same for \(f(g(x))\)

OpenStudy (anonymous):

Ok, I will, even though I dont understand it. Thank you

geerky42 (geerky42):

What don't you understand?

OpenStudy (anonymous):

Why there are two 4s?

geerky42 (geerky42):

Because both \(f(x)\) and \(g(x)\) are equal to \(\dfrac{4}{x}\), right? So you have \(g(\mathbf{\color{red}{x}}) = \dfrac{4}{\mathbf{\color{red}{x}}}\Longrightarrow g(\mathbf{\color{red}{f(x)}}) = \dfrac{4}{\mathbf{\color{red}{\dfrac{4}{x}}}}\) Does that make sense? Color helps?

geerky42 (geerky42):

Basically, you just replace x to f(x) aka 4/x

OpenStudy (anonymous):

Ok thank you.